OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95.

OpenStudy (nincompoop):

do you know what the formula for arithmetic sequence is?

OpenStudy (ankitshaw):

a+ (n-1)d = eqn to find nth term of AP . use it to find the difference and then u easily get 35th term

OpenStudy (nincompoop):

\[a_n=a_1+(n-1)d\] a1 is the verst term d is the common difference n is the number of the term to find

OpenStudy (anonymous):

it would be a 35 = a -7 + (35-(7)) right?

OpenStudy (nincompoop):

you need to know the common difference first

OpenStudy (nincompoop):

-7 is your first term and 35 is the nth term you need to find. there isn't enough data to plug and find out the answer immediately. the common difference, d, must be identified first

OpenStudy (anonymous):

102/17

OpenStudy (anonymous):

6!!!!

OpenStudy (nincompoop):

6 is the common difference?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

try it out

OpenStudy (anonymous):

I know how to plug it into the equation but i have no clue how to solve it.

OpenStudy (nincompoop):

what is your equation?

OpenStudy (nincompoop):

put your equation up and I'll help you solve

OpenStudy (anonymous):

35 = a -7 + (35-(7)) 6

OpenStudy (ilovebmth1234):

so you have to solve for a?

OpenStudy (anonymous):

I have no clue what I have to solve for

OpenStudy (ilovebmth1234):

i think you solve for A in that equation it looks like

OpenStudy (anonymous):

okay. the a has those exponents on the bottom... im not sure what theyre called

OpenStudy (ilovebmth1234):

what you mean? where?

OpenStudy (anonymous):

It's #3

OpenStudy (ilovebmth1234):

197 hun is the answer

OpenStudy (ilovebmth1234):

but let me check it once more to see if im right

OpenStudy (ilovebmth1234):

e little numbers at the bottom of the numbers are telling you wich term they are like -7 is the first term and 95 is the 18th term and ther 35th term would be 197

OpenStudy (anonymous):

did you just multiply?

OpenStudy (ilovebmth1234):

ill write out what i did and how to do it

OpenStudy (anonymous):

oh okay awesome :3

OpenStudy (ilovebmth1234):

n = 35; a1 = -7, d = 6 \[a _{n}=a _{1}+(n-1)d \] \[a _{35}=-7+(35-1)6\] \[a _{35}=197\] The tenth term is 35.

OpenStudy (ilovebmth1234):

make sense?

OpenStudy (ilovebmth1234):

so when you get another problem like this just change the numbers and the equation always stays the same

OpenStudy (anonymous):

:O that makes so much sense! thank you!! :D you are a life saver <3

OpenStudy (ilovebmth1234):

any time lol i just learned about this myself like 3 weeks ago so im glad i could help :D <3 anything you need help with?

OpenStudy (anonymous):

I think that's all I have :) I think I can figure out the rest using your example

OpenStudy (ilovebmth1234):

alright hun im always here to help :)